OpenStudy (saifoo.khan):

An executive buys $2490 worth of presents for the children of her employees. For each girl she gets an art kit costing $33; each boy receives a set of tools costing $29. How many presents of each type did she buy?

OpenStudy (dumbsearch2):

Don't tell me a mod has schoolwork! :O

OpenStudy (saifoo.khan):

Haha. Why? Don't they study?

OpenStudy (dumbsearch2):

I thought of mods as some kind of voodoo god that goes in a cape and flys in the air with their purple helmet of copper and steel.

OpenStudy (mertsj):

We are but like Superman, we often operate undercover.

hartnn (hartnn):

how many total children are there ?

OpenStudy (texaschic101):

33g + 29b = 2490 There is one equation. How many kids in total ?

hartnn (hartnn):

let there be 'g' girls and 'b' boys. 33g + 29 b = 2490 you need more info.

OpenStudy (saifoo.khan):

No idea.

hartnn (hartnn):

do you have choices ?

OpenStudy (saifoo.khan):

My closest guess was 40 for each @hartnn

OpenStudy (tkhunny):

It's harder to get a higher SmartScore if you NEVER ask a question. All Girls $2490 / 33 = 75.45 All Boys $2490 / 29 = 85.86 So, somewhere between 75 and 85 total gifts.

OpenStudy (texaschic101):

would it be g = g - b ? Does that make sense ? Could that be the first equation ? Then sub g - b in for g ?

OpenStudy (saifoo.khan):

I think it will give you the same equation?

OpenStudy (saifoo.khan):

So guys? No idea? :/

OpenStudy (texaschic101):

33g + 29b = 2490 g = g - b 33(g - b) + 29b = 2490 33g - 33b + 29b = 2490 33g - 4b = 2490 33g = 2490 - 4b g = (2490 - 4b)/33 g = 75 - 4b/33 well...this is not working....sorry I don't know

OpenStudy (saifoo.khan):

No problem chic. XD

OpenStudy (texaschic101):

@jim_thompson5910 .....lets see if you can do this one :)

OpenStudy (saifoo.khan):

More screenshots

OpenStudy (dumbsearch2):

I know, @tkhunny. I was just joking ;P

OpenStudy (dumbsearch2):

haha @Mertsj

OpenStudy (saifoo.khan):

Saw it already. Unfortunately they haven't answered the question. :/

OpenStudy (dumbsearch2):

Honestly I thought that the mods were just lazy guys just sitting back who rarely even use the site ;P

OpenStudy (saifoo.khan):

Lol. There are all types of people and all types of mods too. ;)

OpenStudy (dumbsearch2):

Thats why the moderation is so bad, esp. in areas like the Music section/chat. I'm still kinda mad about this. Even in the math chat things get CoC-violated.

OpenStudy (mertsj):

@dumbsearch2 This is America...you have a right to be wrong.

OpenStudy (saifoo.khan):

^

OpenStudy (dumbsearch2):

W-O-W. Then what's the point of mods ?

OpenStudy (dumbsearch2):

Today, with my VERY OWN EYES, I witnessed tons of inappropriate chat stuff.

OpenStudy (saifoo.khan):

To keep the wrong thing in limits or on track? Maybe?

OpenStudy (saifoo.khan):

Your eyes were deceiving you. XD

OpenStudy (dumbsearch2):

Oh. My screenshot tool doesn't deceive me xD

OpenStudy (dumbsearch2):

its alright. this won't change. it's been like this for months. I'll stop my ramblings. Go back to what you were doing (:

hartnn (hartnn):

oh, so this is the problem of diophantine equations... stop arguing ?

OpenStudy (saifoo.khan):

But your eyes do. :P People are thinking about it i guess..

OpenStudy (anonymous):

\[x = \text{number of girls}=\text{number of knits}\\ y = \text{number of boys}=\text{number of tool kits}\\ 33x+29y=2490\\ x+y = N\\ 33x+29(N-x)=2490\\ 4x=2490-29N\\ x=622.5-29{N\over 4} \] hint 1: N must be an odd integer

OpenStudy (anonymous):

i.e., there are in all a odd number of children

OpenStudy (anonymous):

now, \[622.5\ge 29{2n+1\over 4}\]

OpenStudy (saifoo.khan):